# Learn from Global Correlations: Enhancing Evolutionary Algorithm via Spectral GNN

### v4

@@ Introduction
Graph Neural Networks (GNNs) have been widely applied in fields such as social network analysis, recommendation systems, and protein structure prediction because of their capacity to effectively model complex relational structures using graph representations(Kipf and Welling 2016; Veličković et al. 2017; Xu et al. 2018; Wang et al. 2019; Hamilton, Ying, and Leskovec 2017). In GNNs, dependencies between nodes are captured by propagating and aggregating information across the network, thereby enabling adaptive learning and optimization based on node interactions. Similarly, biological evolution involves genetic interactions characterized by variation, recombination, and natural selection, which are processes that collectively facilitate the evolutionary optimization of genotypes toward greater fitness(Wright et al. 1932; Darwin 2023; Brady 1985). Given these parallels, we propose that GNNs can be understood through an evolutionary perspective, where the propagation of information among nodes mirrors the diversity-exploring mechanisms of crossover and mutation, as well as the adaptive optimization mechanism of directional selection in evolution.
@@
Evolutionary Algorithms. EAs are a family of metaheuristic optimization techniques inspired by natural evolutionary processes, incorporating operators such as natural selection and mutation. GA is one of the earliest and most representative EAs, and is particularly effective in solving discrete and multimodal optimization problems(Holland 1992). In recent years, DE(Storn and Price 1997; Opara and Arabas 2019; Pant et al. 2020) and Evolution Strategy (ES)(Beyer and Schwefel 2002; Lange, Tian, and Tang 2024; Beyer and Arnold 2001) have emerged as prominent research domains. DE generates candidate solutions through differential mutation and selects the fittest among them. Numerous variants have been developed, including Adaptive Differential Evolution with Optional External Archive (JADE) (Zhang and Sanderson 2009), which incorporates external archives and adaptive parameter updates; Success-History Based Parameter Adaptation for Differential Evolution (SHADE) (Tanabe and Fukunaga 2013), which utilizes historically successful parameter settings to refine control parameters; and RL-SHADE (Zhang et al. 2022), which further enhances DE performance by integrating natural evolution strategies for more effective parameter optimization. Among ES methods, Covariance Matrix Adaptation Evolution Strategy (CMA-ES) is one of the most widely recognized algorithms (Hansen, Müller, and Koumoutsakos 2003), adjusting the covariance matrix to effectively explore the solution space. Several variants have been developed, including OpenES for large-scale parallel computation (Salimans et al. 2017); Parameter-Exploring Policy Gradients (PEPG) for estimating gradients in the parameter space (Sehnke et al. 2010);Mixture Model based Evolution Strategy (MMES) (He, Zheng, and Zhou 2020), which accelerates CMA-ES in large-scale optimization tasks; and SDAES (He et al. 2019), which reduces both time and space complexity in high-dimensional settings. From GA to DE, ES, and their variants, the design of evolutionary algorithms has consistently centered on balancing exploration and exploitation. Exploration introduces diversity to expand the search space and prevent premature convergence to local optima, while exploitation refines promising solutions to gradually approach the global optimum(Eiben and Schippers 1998).
@@
Graph Neural Networks. GNNs process graph-structured data from two principal perspectives: spatial and spectral(Chen et al. 2023; Muhammet et al. 2020; Wang and Zhang 2022). Spatial GNNs are more widely known, with the Graph Convolutional Network (GCN) (Kipf and Welling 2016) being a representative example, which aggregates information from neighboring nodes to learn local structures. The Graph Attention Network (GAT) (Veličković et al. 2017) further enhances GCNs by incorporating self-attention mechanisms, while GraphSAGE generates node embeddings by sampling and aggregating features from local neighborhoods, thereby enabling generalization to previously unseen nodes (Hamilton, Ying, and Leskovec 2017). In contrast, spectral GNNs rely on the graph Laplacian and Fourier transform for convolution operations, thus learning from global structural patterns. Here we review representative spectral methods. ChebyNet introduces Chebyshev polynomials to construct efficient spectral filters, thereby improving scalability (Defferrard, Bresson, and Vandergheynst 2016). BernNet learns spectral filters through Bernstein polynomial approximations, which provide flexibility in filter design and help overcome the limitations of predefined or unconstrained filters (He et al. 2021). JacobiConv increases expressive power by employing an orthogonal Jacobi basis, yielding superior performance over other spectral GNNs across both synthetic and real-world datasets (Wang and Zhang 2022). In brief, spectral GNNs derive filtering expressions that modulate both the low-frequency and high-frequency components of the original graph signals. This process aligns with evolutionary dynamics: low-frequency components represent the stable core of the graph signal, while high-frequency components capture variations among nodes, thereby capturing finer details.
@@
Notations. Given a set of nodes $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, an undirected graph can be represented in matrix form as $\mathcal{G}=(\mathbf{X},\mathbf{A})$, where the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the pairwise interactions between nodes.The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$, and the normalized Laplacian matrix is presented as $\widehat{\mathbf{L}}\ =\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. This matrix can also be expanded in the spectral form as $\widehat{\mathbf{L}}\ =\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ is a diagonal matrix of the corresponding eigenvalues, which satisfy $\lambda_{k}\in[0,2]$. The graph Fourier transform is defined as $\widetilde{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$, with inverse transform $\mathbf{X}=\mathbf{U}\widetilde{\mathbf{X}}$. The component at frequency $\lambda_{k}$ is given by $\widetilde{\mathbf{X}}_{\lambda_{k}}=\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}$, where $\mathbf{U}_{k}$ is the eigenvector corresponding to $\lambda_{k}$. The spectral filter is defined as $g(\mathbf{\Lambda})=\mathrm{diag}\bigl{(}g(\lambda_{1}),\dots,g(\lambda_{N})\bigr{)}$ and is generally parameterized by a $k$th-order polynomial, $g(\lambda)=\sum_{k=0}^{K}\alpha_{k}\lambda^{k}$, which lifts to the vertex domain through $g(\widehat{\mathbf{L}})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}$. Problem formulation. In the proposed GNE framework, the evolution of the population can be regarded as a frequency filtering process. By controlling the frequency components, the balance between exploration and exploitation is achieved. Specifically, the GNE population-update procedure is based on a polynomial filter and can be defined as follows:
@@
where $\mathbf{U}_{ki}$ and $\mathbf{U}_{kj}$ denote the components of individuals $i$ and $j$ in the eigenvector $\mathbf{U}_{t}$, and $d_{i}$ and $d_{j}$ represent their degrees. The eigenvalues $\lambda_{k}$ have a physical meaning as frequencies and quantify the variability along $\mathbf{U}_{k}$. Higher values of $\lambda_{k}$ indicate greater variability, with transitions from high to low frequencies reflecting a shift from diverse traits to consistent traits, which aligns with maintaining population diversity in EAs to fully explore the solution space while preserving consistency for fine exploration around the optimal solution. Figure 3 provides a vivid example visualizing this relationship.
@@
2:Output: Optimal solution $\mathbf{x}_{\mathrm{best}}$ and objective value $f_{\mathrm{best}}$.
3: Initialize population $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ with random solutions.
@@
8:  Encode population with $\varphi$: $\mathbf{X}\leftarrow\varphi(\mathbf{X})$.
9:  Compute similarity matrix $\mathbf{A}$ from $\mathbf{X}$.
@@
11:  Compute normalized Laplacian $\widehat{\mathbf{L}}\leftarrow\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$.
@@
where the term $\widetilde{\mathbf{X}}_{\lambda_{k}}$ represents the population’s component corresponding to $\lambda_{k}$ , while $g(\lambda_{k})$ is the weight applied to this component. Increasing the weight of high-frequency components boosts exploration, while increasing the weight of low-frequency components strengthens exploitation. The updated population $\hat{\mathbf{X}}$ is obtained by integrating different frequency components. GNNs typically employ multilayer perceptron (MLP) encoders $\psi(\cdot)$ and $\phi(\cdot)$ to transform and map features both before and after filtering, thereby further optimizing node representations. Similarly, GNE treats $\psi(\cdot)$ and $\phi(\cdot)$ as sampling functions that guide the population to generate new individuals toward optimizing overall population fitness. Here, $\psi(\cdot)$ is an identity transformation, while $\phi(\cdot)$ is a Gaussian distribution centered at the current best solution $\mathbf{x}_{\text{best}}$ with variance $0.0005\cdot\exp\left(-16\cdot\frac{t-T/2}{T/2}\right)$, where $t$ is the current iteration and $T$ is the maximum number of iterations.To summarize, the overall update formula for GNE can be expressed as follows:

### v5

@@ Introduction
Graph Neural Networks (GNNs) have been widely applied in fields such as social network analysis, recommendation systems, and protein structure prediction because of their capacity to effectively model complex relational structures using graph representations(Kipf 2016; Petar et al. 2018; Xu et al. 2018; Liu et al. 2025; Lu et al. 2025; Zhou et al. 2025; Cui et al. 2020). In GNNs, dependencies between nodes are captured by propagating and aggregating information across the network(Yu et al. 2025, 2024; Guan et al. 2025b, c, a), thereby enabling adaptive learning and optimization based on node interactions. Similarly, biological evolution involves genetic interactions characterized by variation, recombination, and natural selection, which are processes that collectively facilitate the evolutionary optimization of genotypes toward greater fitness(Wright et al. 1932; Darwin 2023; Brady 1985). Given these parallels, we propose that GNNs can be understood through an evolutionary perspective, where the propagation of information among nodes mirrors the diversity-exploring mechanisms of crossover and mutation, as well as the adaptive optimization mechanism of directional selection in evolution.
@@
Evolutionary Algorithms. EAs are a family of metaheuristic optimization techniques inspired by natural evolutionary processes, incorporating operators such as natural selection and mutation(Ouyang et al. 2025b, a). GA is one of the earliest and most representative EAs, and is particularly effective in solving discrete and multimodal optimization problems(Holland 1992). In recent years, DE(Storn and Price 1997; Opara and Arabas 2019; Pant et al. 2020) and Evolution Strategy (ES)(Beyer and Schwefel 2002; Lange, Tian, and Tang 2024; Beyer and Arnold 2001) have emerged as prominent research domains. DE generates candidate solutions through differential mutation and selects the fittest among them. Numerous variants have been developed, including Adaptive Differential Evolution with Optional External Archive (JADE) (Zhang and Sanderson 2009), which incorporates external archives and adaptive parameter updates; Success-History Based Parameter Adaptation for Differential Evolution (SHADE) (Tanabe and Fukunaga 2013), which utilizes historically successful parameter settings to refine control parameters; and RL-SHADE (Zhang et al. 2022), which further enhances DE performance by integrating natural evolution strategies for more effective parameter optimization. Among ES methods, Covariance Matrix Adaptation Evolution Strategy (CMA-ES) is one of the most widely recognized algorithms (Hansen, Müller, and Koumoutsakos 2003), adjusting the covariance matrix to effectively explore the solution space. Several variants have been developed, including OpenES for large-scale parallel computation (Salimans et al. 2017); Parameter-Exploring Policy Gradients (PEPG) for estimating gradients in the parameter space (Sehnke et al. 2010);Mixture Model based Evolution Strategy (MMES) (He, Zheng, and Zhou 2020), which accelerates CMA-ES in large-scale optimization tasks; and SDAES (He et al. 2019), which reduces both time and space complexity in high-dimensional settings. From GA to DE, ES, and their variants, the design of evolutionary algorithms has consistently centered on balancing exploration and exploitation. Exploration introduces diversity to expand the search space and prevent premature convergence to local optima, while exploitation refines promising solutions to gradually approach the global optimum(Eiben and Schippers 1998).
@@
Graph Neural Networks. GNNs process graph-structured data from two principal perspectives: spatial and spectral(Chen et al. 2023; Muhammet et al. 2020; Wang and Zhang 2022; Meng et al. 2024b, a; Yang et al. 2021, 2022). Spatial GNNs are more widely known, with the Graph Convolutional Network (GCN) (Kipf 2016) being a representative example, which aggregates information from neighboring nodes to learn local structures. The Graph Attention Network (GAT) (Petar et al. 2018) further enhances GCNs by incorporating self-attention mechanisms, while GraphSAGE generates node embeddings by sampling and aggregating features from local neighborhoods, thereby enabling generalization to previously unseen nodes (Hamilton, Ying, and Leskovec 2017). In contrast, spectral GNNs rely on the graph Laplacian and Fourier transform for convolution operations, thus learning from global structural patterns. Here we review representative spectral methods. ChebyNet introduces Chebyshev polynomials to construct efficient spectral filters, thereby improving scalability (Defferrard, Bresson, and Vandergheynst 2016). BernNet learns spectral filters through Bernstein polynomial approximations, which provide flexibility in filter design and help overcome the limitations of predefined or unconstrained filters (He et al. 2021). JacobiConv increases expressive power by employing an orthogonal Jacobi basis, yielding superior performance over other spectral GNNs across both synthetic and real-world datasets (Wang and Zhang 2022). In brief, spectral GNNs derive filtering expressions that modulate both the low-frequency and high-frequency components of the original graph signals. This process aligns with evolutionary dynamics: low-frequency components represent the stable core of the graph signal, while high-frequency components capture variations among nodes, thereby capturing finer details.
@@
Notations. Given a set of nodes $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, an undirected graph can be represented in matrix form as $\mathcal{G}=(\mathbf{X},\mathbf{A})$, where the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ encodes the pairwise interactions between nodes.The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{D}_{ii}=\sum_{j}\mathbf{A}_{ij}$, and the normalized Laplacian matrix is presented as $\widehat{\mathbf{L}}\ =\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. This matrix can also be expanded in the spectral form as $\widehat{\mathbf{L}}\ =\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$ is a diagonal matrix of the corresponding eigenvalues, which satisfy $\lambda_{k}\in[0,2]$. The graph Fourier transform is defined as $\widetilde{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$, with inverse transform $\mathbf{X}=\mathbf{U}\widetilde{\mathbf{X}}$. The component at frequency $\lambda_{k}$ is given by $\widetilde{\mathbf{X}}_{\lambda_{k}}=\mathbf{U}_{k}\mathbf{U}_{k}^{\top}\mathbf{X}$, where $\mathbf{U}_{k}$ is the eigenvector corresponding to $\lambda_{k}$. The spectral filter is defined as $g(\mathbf{\Lambda})=\mathrm{diag}\bigl(g(\lambda_{1}),\dots,g(\lambda_{N})\bigr)$ and is generally parameterized by a $k$th-order polynomial, $g(\lambda)=\sum_{k=0}^{K}\alpha_{k}\lambda^{k}$, which lifts to the vertex domain through $g(\widehat{\mathbf{L}})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}$.
Problem formulation. In the proposed GNE framework, the evolution of the population can be regarded as a frequency filtering process. By controlling the frequency components, the balance between exploration and exploitation is achieved. Specifically, the GNE population-update procedure is based on a polynomial filter and can be defined as follows:
@@
2:Output: Optimal solution $\mathbf{x}_{\mathrm{best}}$, objective value $f_{\mathrm{best}}$.
3: Initialize population $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ randomly.
@@
8:  Encode population with $\psi(\cdot)$: $\mathbf{X}\leftarrow\psi(\cdot)(\mathbf{X})$.
9:  Compute adjacency matrix $\mathbf{A}$ from $\mathbf{X}$.
@@
11:  Compute normalized $\widehat{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$.
@@
where $\mathbf{U}_{ki}$ and $\mathbf{U}_{kj}$ denote the components of individuals $i$ and $j$ in the eigenvector $\mathbf{U}_{k}$, and $d_{i}$ and $d_{j}$ represent their degrees. The eigenvalues $\lambda_{k}$ have a physical meaning as frequencies and quantify the variability along $\mathbf{U}_{k}$. Higher values of $\lambda_{k}$ indicate greater variability, with transitions from high to low frequencies reflecting a shift from diverse traits to consistent traits, which aligns with maintaining population diversity in EAs to fully explore the solution space while preserving consistency for fine exploration around the optimal solution. Figure 3 provides a vivid example visualizing this relationship.
@@
where the term $\widetilde{\mathbf{X}}_{\lambda_{k}}$ represents the population’s component corresponding to $\lambda_{k}$, while $g(\lambda_{k})$ is the weight applied to this component. Increasing the weight of high-frequency components boosts exploration, while increasing the weight of low-frequency components strengthens exploitation. The updated population $\hat{\mathbf{X}}$ is obtained by integrating different frequency components. GNNs typically employ multilayer perceptron (MLP) encoders $\psi(\cdot)$ and $\phi(\cdot)$ to transform and map features both before and after filtering, thereby further optimizing node representations. Similarly, GNE treats $\psi(\cdot)$ and $\phi(\cdot)$ as sampling functions that guide the population to generate new individuals toward optimizing overall population fitness. Here, $\psi(\cdot)$ is an identity transformation, while $\phi(\cdot)$ is a Gaussian distribution centered around the elite individuals of the population, with a standard deviation $\sigma_{t}$ that adaptively varies with the iteration count $t$.To summarize, the overall update formula for GNE can be expressed as follows:
@@ Conclusion
Acknowledgments
This work was supported by the Fundamental Research Funds for the Central Universities (No. N25XQD062) and Open Research Projects of the Key Laboratory of Blockchain Technology and Data Security of the Ministry of Industry and Information Technology for the year 2025 (No. KT20250015). Corresponding authors: Puning Zhao and Dayu Hu.